### v1

@@ 2 $S$-Primary Decomposition in $S$-Noetherian Ring
For any $x\in R$, we have $(I:x)=(\bigcap\limits_{i=1}^{n}Q_{i}:x)=\bigcap\limits_{i=1}^{n}(Q_{i}:x)$, and so $rad(I:x)=\bigcap\limits_{i=1}^{n}rad(Q_{i}:x)$. First we show that $rad(I:x)$ satisfies $sP_{i}\subseteq s^{\prime}(rad(I:x))\subseteq P_{i}$ for some $s,s^{\prime}\in S$ is $S$-prime. For this, let $\alpha\beta\in rad(I:x)$ such that $t\beta\notin rad(I:x)$ for all $t\in S$. This implies $s^{\prime}\alpha\beta\in s^{\prime}rad(I:x)\subseteq P_{i}$. Then there exists $s^{\prime\prime}\in S$ such that either $s^{\prime\prime}\alpha\in P_{i}$ or $s^{\prime}s^{\prime\prime}\beta\in P_{i}$ since $P_{i}$ is $S$-prime. If $s^{\prime}s^{\prime\prime}\beta\in P_{i}$, then $ss^{\prime}s^{\prime\prime}\beta\in sP_{i}\subseteq s^{\prime}(rad(I:x))\subseteq rad(I:x)$, a contradiction since $t\beta\notin rad(I:x)$ for all $t\in S$. Hence $s^{\prime\prime}\alpha\in P_{i}$, and so $ss^{\prime\prime}\alpha\in sP_{i}\subseteq s^{\prime}rad(I:x)\subseteq rad(I:x)$, i.e., $s^{\prime\prime\prime}\alpha\in rad(I:x)$, where $s^{\prime\prime\prime}=ss^{\prime\prime}\in S$. Thus $rad(I:x)$ is an $S$-prime ideal. Now by Lemma 14, there exist $t^{\prime}\in S$ and $j\in\{1,2,\ldots,n\}$ such that $t^{\prime}(rad(Q_{j}:x))\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$. Choose $x$ such that $tx\notin Q_{j}$ for all $t\in S$. Then by Lemma 15, there exists $t^{\prime\prime}\in S$ such that $t^{\prime\prime}(rad(Q_{j}:x))\subseteq rad(Q_{j})=P_{j}\subseteq rad(Q_{j}:x)$. This implies that $t^{\prime}t^{\prime\prime}(rad(Q_{j}:x))\subseteq t^{\prime}P_{j}\subseteq t^{\prime}(rad(Q_{j}:x))\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$. Thus we have $t^{\prime}P_{j}\subseteq rad(I:x)\subseteq rad(Q_{j}:x)$, and so $t^{\prime\prime}t^{\prime}P_{j}\subseteq t^{\prime\prime}(rad(I:x))\subseteq t^{\prime\prime}(rad(Q_{j}:x))\subseteq rad(Q_{j})=P_{j}$. Put $s=t^{\prime\prime}t^{\prime}$ and $s^{\prime}=t^{\prime\prime}$. Then we have $sP_{j}\subseteq s^{\prime}(rad(I:x))\subseteq P_{j}$ for some $j$. Conversely, since decomposition is minimal, so for each $k$, $\bigcap\limits_{i\neq k}^{n}Q_{i}\nsubseteq Q_{k}$. This implies that there exists $x_{k}\in(\bigcap\limits_{i\neq k}^{n}Q_{i})\setminus Q_{k}$ for each $k$. Consequently, for each $k$, we have $(I:x_{k})=\bigcap\limits_{i=1}^{n}(Q_{i}:x_{k})=(Q_{k}:x_{k})$ since $(Q_{i}:x_{k})=R$ for all $i\neq k$ by [7, Lemma 4.4]. Notice that $tx_{k}\notin Q_{k}$ for every $t\in S$, for if $tx_{k}\in Q_{k}$ for some $t\in S$, then $t\in(Q_{k}:x_{k})=(I:x_{k})\subseteq rad(I:x_{k})$, which is not possible since $rad(I:x_{k})$ is $S$-prime. So by Lemma 15, for each $k$, there exists $s_{k}\in S$ such that $s_{k}(rad(Q_{k}:x_{k}))\subseteq rad(Q_{k})=P_{k}$. Also $Q_{k}\subseteq(Q_{k}:x_{k})=(I:x_{k})$, so $P_{k}=rad(Q_{k})\subseteq rad(I:x_{k})$. Consequently, for each $k$, $s_{k}P_{k}\subseteq s_{k}(rad(I:x_{k}))=s_{k}(rad(Q_{k}:x_{k}))\subseteq P_{k}$. Put $s=s_{1}s_{2}\ldots s_{n}$. Then $sP_{k}\subseteq s(rad(I:x_{k}))\subseteq P_{k}$ for all $k=1,2,\ldots,n$. This completes the proof. ∎
For any ideal $I$ of a ring $R$ and any multiplicative set $S\subseteq R$, the contraction of ideal $S^{-1}I$ is denoted by $S(I)=\{a\in R\hskip 5.69046pt|\hskip 5.69046pt\dfrac{a}{1}\in S^{-1}I\}$. Notice that $I\subseteq S(I)$. To prove the second uniqueness theorem for the $S$-Noetherian rings, we need the following results:
Lemma 17.
Let $R$ be a ring, $S\subseteq R$ be a multiplicative set, and $Q$ be an $S$-$P$-primary ideal of $R$. If $S\cap P=\emptyset$, then $S^{-1}Q$ is $S^{-1}P$-primary and its contraction in $R$ satisfies $tS(Q)\subseteq Q\subseteq S(Q)$ for some $t\in S$.
Proof.
Clearly $S^{-1}Q\neq S^{-1}R$, for $S\cap P=\emptyset$ and $P=rad(Q)$. Let $x\in S(Q)$. Then $\dfrac{x}{1}=\dfrac{a}{s}$ for some $a\in Q$ and $s\in S$. Consequently, there exists $s^{\prime}\in S$ such that $(xs-a)s^{\prime}=0$; whence $ss^{\prime}x=s^{\prime}a\in Q$. Put $t=ss^{\prime}$. Then $tx\in Q$, and so $tS(Q)\subseteq Q\subseteq S(Q)$. Now since $Q$ is $S$-primary, by [16, Proposition 2.6], $S^{-1}Q$ is a primary ideal of $S^{-1}R$ and $rad(S^{-1}Q)=S^{-1}(rad(Q))=S^{-1}P$. Also since $P$ is an $S$-prime ideal of $R$, by [2, Remark $1$], $S^{-1}P$ is a prime ideal of $S^{-1}R$. Thus $S^{-1}Q$ is an $S^{-1}P$-primary ideal of $S^{-1}R$.
Lemma 18.
Let $R$ be an $S$-Noetherian ring, and $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $S$-$P_{i}$-primary. If $S\cap\left(\bigcap\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ for some $m$. Then $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$ and $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S(I)$ for some $t\in S$.
Proof.
By [7, Proposition 3.11(v)], $S^{-1}I=\bigcap\limits_{i=1}^{n}S^{-1}Q_{i}$. Since $S\cap Q_{i}\neq\emptyset$ for $i=m+1,m+2,\ldots,n$, so by [7, Proposition 4.8(i)], $S^{-1}Q_{i}=S^{-1}R$ for $i=m+1,m+2,\ldots,n$. Consequently, $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$. Also by Lemma 17, $S^{-1}Q_{i}$ is $S^{-1}P_{i}$-primary for $i=1,2,\ldots,m$. Since $P_{i}$ are distinct, $S^{-1}P_{i}$ are also distinct, and so $S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$ is a minimal $S$-primary decomposition. Next, let $x\in S(I)$. Then $\dfrac{x}{1}\in S^{-1}I=\bigcap\limits_{i=1}^{m}S^{-1}Q_{i}$. Write $\dfrac{x}{1}=\dfrac{a_{i}}{s_{i}}$ for some $a_{i}\in Q_{i}$ and $s_{i}\in S$. This implies that $(xs_{i}-a_{i})s=0$ for some $s\in S$; whence $s_{i}sx=sa_{i}\in Q_{i}$ for $i=1,2,\ldots,m$. Put $t=ss_{1}s_{2}\ldots s_{m}$. Then $tx\in\bigcap\limits_{i=1}^{m}Q_{i}$, and hence $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}$. Also by [7, Proposition 1.17(i)], $\bigcap\limits_{i=1}^{m}Q_{i}\subseteq\bigcap\limits_{i=1}^{m}S(Q_{i})=S(I)$. Hence $tS(I)\subseteq\bigcap\limits_{i=1}^{m}Q_{i}\subseteq S(I)$. ∎
For an ideal $I$ of $R$, let $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition, where $Q_{i}$ is $S$-$P_{i}$-primary. Then the $S$-prime ideals $P_{i}$ are said to belong to $I$. Also we say that $P_{i}$ ($1\leq i\leq n$) is isolated $S$-prime if $sP_{j}\nsubseteq P_{i}$ for all $s\in S$ and for all $j\neq i$. Other $S$-prime ideals are called embedded $S$-prime ideals.
Theorem 19.
($S$-version of 2nd uniqueness theorem). Let $R$ be an $S$-Noetherian ring and $I=\bigcap\limits_{i=1}^{n}Q_{i}$ be a minimal $S$-primary decomposition of $I$, where $Q_{i}$ is $S$-$P_{i}$-primary. If $\{P_{{1}},\ldots,P_{{m}}\}$ is a set of isolated $S$-prime ideals of $I$ for some $m$ $(1\leq m\leq n)$, then $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition.
Proof.
@@
Since each $P_{i}$ is $S$-prime, so by [2, Proposition 1], there exist $s_{i}\in S$ such that $(P_{i}:s_{i})$ is a prime ideal for $i=1,2,\ldots,m$. Consider a multiplicative set $S^{\prime}=R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$ of $R$. Evidently, $S^{\prime}\cap P_{i}=\emptyset$ for $i=1,2,\ldots m$ since $P_{i}\subseteq(P_{i}:s_{i})$. We claim that $S^{\prime}\cap P_{k}\neq\emptyset$ for $k=m+1,m+2,\ldots,n$. On contary $S^{\prime}\cap P_{k}=\emptyset$ for some $k$. This implies that $P_{k}\subseteq\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$, so by [7, Proposition 1.11(i)], $P_{k}\subseteq(P_{j}:s_{j})$ for some $j$ $(1\leq j\leq m)$. Consequently, $s_{j}P_{k}\subseteq P_{j}$, a contradiction since $P_{j}$ is isolated $S$-prime. This concludes that $S^{\prime}\cap\left(\bigcap\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S^{\prime}\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ since $P_{i}=rad(Q_{i})$ for all $i=1,2,\ldots,n.$ Now since each $P_{i}$ is $S$-prime, $S\cap P_{i}=\emptyset$ for all $i=1,2,\ldots,m$. This implies that $S\cap(P_{i}:s_{i})=\emptyset$ for all $i=1,2,\ldots,m$. For, if $S\cap(P_{i}:s_{i})\neq\emptyset$ for some $i$, then there exits $s^{\prime}\in S$ such that $s^{\prime}s_{i}\in P_{i}$, a contradiction. Consequently, $S\cap(\bigcup\limits_{i=1}^{m}(P_{i}:s_{i}))=\emptyset$, and so $S\subseteq R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})=S^{\prime}$. Then by [5, Proposition 2(d)], $R$ is $S^{\prime}$-Noetherian. Also since each $Q_{i}$ is $S$-primary and $S^{\prime}\cap Q_{i}=\emptyset$ for all $i$ $(1\leq i\leq m)$, so $Q_{i}$ is $S^{\prime}$-primary for all $i=1,2,\ldots,m.$ Thus by Lemma 18, we have $tS(I)\subseteq\bigcap\limits_{j=1}^{m}Q_{j}\subseteq S(I)$ for some $t\in S$. Hence $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition. ∎

### v2

@@ 2 $S$-Primary Decomposition in $S$-Noetherian Ring
Since each $P_{i}$ is $S$-prime, by [2, Proposition 1], there exists $s_{i}\in S$ such that $(P_{i}:s_{i})$ is a prime ideal for $i=1,2,\ldots,m$. Consider the multiplicative set $S^{\prime}=R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$ of $R$. Evidently, $S^{\prime}\cap P_{i}=\emptyset$ for $i=1,2,\ldots m$ since $P_{i}\subseteq(P_{i}:s_{i})$. We claim that $S^{\prime}\cap P_{k}\neq\emptyset$ for $k=m+1,m+2,\ldots,n$. On the contrary $S^{\prime}\cap P_{k}=\emptyset$ for some $k$. This implies that $P_{k}\subseteq\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})$, by [6, Proposition 1.11(i)], $P_{k}\subseteq(P_{j}:s_{j})$ for some $j$ $(1\leq j\leq m)$. Consequently, $s_{j}P_{k}\subseteq P_{j}$, a contradiction since $P_{j}$ is isolated $S$-prime. This concludes that $S^{\prime}\cap\left(\bigcup\limits_{i=1}^{m}Q_{i}\right)=\emptyset$ and $S^{\prime}\cap\left(\bigcap\limits_{i=m+1}^{n}Q_{i}\right)\neq\emptyset$ since $P_{i}=rad(Q_{i})$ for all $i=1,2,\ldots,n.$ Now since each $P_{i}$ is $S$-prime, $S\cap P_{i}=\emptyset$ for all $i=1,2,\ldots,m$. This implies that $S\cap(P_{i}:s_{i})=\emptyset$ for all $i=1,2,\ldots,m$. For, if $S\cap(P_{i}:s_{i})\neq\emptyset$ for some $i$, then there exists $s^{\prime}\in S$ such that $s^{\prime}s_{i}\in P_{i}$, a contradiction. Consequently, $S\cap(\bigcup\limits_{i=1}^{m}(P_{i}:s_{i}))=\emptyset$, and so $S\subseteq R-\bigcup\limits_{i=1}^{m}(P_{i}:s_{i})=S^{\prime}$. Also since each $Q_{i}$ is $S$-primary and $S^{\prime}\cap Q_{i}=\emptyset$ for all $i$ $(1\leq i\leq m)$, $Q_{i}$ is $S^{\prime}$-primary for all $i=1,2,\ldots,m.$ Thus by Lemma 18, we have $tS^{\prime}(I)\subseteq\bigcap\limits_{j=1}^{m}Q_{j}\subseteq S^{\prime}(I)$ for some $t\in S$. From above it is clear that $Q_{{1}}\cap\cdots\cap Q_{{m}}$ depends only $I$, and hence $Q_{{1}}\cap\cdots\cap Q_{{m}}$ is independent of $S$-primary decomposition. ∎
Acknowledgement: Authors sincerely thank the anonymous referee for thorough review and very useful comments and suggestions that helped to improve this paper.